The given points are vertices of a right triangle. Find the length of the hypotenuse of each triangle. (-4,-7) (0,6) (7,5)
Answer by Boreal(15235)

Distance formula between first and second, first and third, and second and third works.
first and second is sqrt(16+169)=sqrt (185)
second and third is sqrt (50)
first and third is sqrt (121+144)=sqrt (265)
This is not a right triangle. This is proven by graphing as well, since no two lines are perpendicular, which they must be.
I reworked the slopes of all the lines, and they are 13/4, -1,7, and 12/11. I can't find a negative reciprocal.